Lewis Company’s standard labor cost of producing one unit of Product DD is 3.20 hours at the rate of $13.80 per hour. During August, 41,200 hours of labor are incurred at a cost of $14.00 per hour to produce 12,600 units of Product DD.

(a)

Compute the total labor variance.

Total labor variance $

UnfavorableNeither favorable nor unfavorableFavorable


(b)

Compute the labor price and quantity variances.

Labor price variance $

UnfavorableFavorableNeither favorable nor unfavorable

Labor quantity variance $

UnfavorableNeither favorable nor unfavorableFavorable


(c)

Compute the labor price and quantity variances, assuming the standard is 3.60 hours of direct labor at $14.10 per hour.

Labor price variance $

FavorableNeither favorable nor unfavorableUnfavorable

Labor quantity variance $

Neither favorable nor unfavorableFavorableUnfavorable

Click if you would like to Show Work for this question:

Open Show Work
Answer #1

(a)
Total labor variance = Actual Labor cost - Standard Labor cost
= (41200 x $14) - (12600 x 3.20 x $13.80) = $20384 (U)

(b)
Labor Price Variance = (Actual rate - Standard Rate) x Actual hours
= ($14 - $13.80) x 41200 = $8240 (U)

Labor Quantity Variance = (Actual Hours - Standard Hours) x Standard Rate
= (41200 - 40320) x $13.80 = $12144 (U)
Standard Hours = 12600 x 3.2 = 40320 hours

(c)
Labor Price Variance = (Actual rate - Standard Rate) x Actual hours
= ($14 - $14.10) x 41200 = $4120 (F)

Labor Quantity Variance = (Actual Hours - Standard Hours) x Standard Rate
= (41200 - 45360) x $14.10 = $58656 (F)
Standard Hours = 12600 x 3.6 = 45360 hours